S Band High Gregorian Dome Receiver Log

as of Friday, 03-Dec-2021 06:21:07 AST

Go to Most Recent Entry

ANS anolla@naic.edu 3/22/2002 System installed , its rf output is at position 8. The Tsys for this system is 28 k.
ANS anolla@naic.edu 5/14/2002 System down due to cryo comp failure. Flappers apparently broken. Tommorow we will swith cryo comp, from ED 013 (bad one) to spare one ED 010.
ANS anolla@naic.edu 5/15/02 Replaced compressor ED013 with ED010, started cooling the receiver at 11:15am. Vacuum OK.
ANS anolla@naic.edu 5/16/2002 Found that the system vacuum was bad (sweating Dewar).So open the valve again and started cool down .
LSWlwray@naic.eduFriday, September 6, 2002 17:28:59Automatic shutter installed this week. It works in the same manner as the SBW and CB shutters. use "shutter" command to control. Inside of shutter's cover has absorber on it.
ansanolla@naic.eduMonday, September 9, 2002 15:40:19The shutter was to be removed because it was making that the others shutters as CB & SBW doesn't work.
LSWlwray@naic.eduTuesday, September 24, 2002 16:20:07Shutter has been fixed and re-installed. It is operational as of this afternoon.
LSWlwray@naic.eduWednesday, October 2, 2002 13:48:18Post amp chassis is down in the lab. Odd Tsys values for Ch B (negative). Receiver is down.
ansanolla@naic.eduFriday, October 11, 2002 13:25:19Channel B amp 3rd stage shorted out.
LSWlwray@naic.eduThursday, October 24, 2002 08:27:59We warmed up the receiver on Oct 14 (?) in order to make changes inside the dewar. (Check on amplifier)
LSWlwray@naic.eduWednesday, November 20, 2002 17:04:03Due to compressor problems, this receiver is down as of Nov 20, expected to be back on line in one week.
ansanolla@naic edu.Thursday, December 5, 2002 15:32:24cryo system off (cryopump in order to replace it), syst cryo temp too high
ansanolla@naic edu.Friday, December 6, 2002 13:00:25Cryopump replaced with ED 013 , it was ED 007. Been cooled down.
ansanolla@naic edu.Wednesday, December 11, 2002 14:44:22System turned off , cryo system is not cooling the Dewar enough , termal short suspected . Dewar is condensing water outside of it.
ansanolla@naic edu.Tuesday, December 17, 2002 13:09:26Swaping cryocomp it fall and broke cryo lines , so we started pumping cryo lines and pump.
ans anolla@naic edu.Wednesday, December 18, 2002 08:52:10System was cooled down , cryocomp is outside its shack , so need to put it back on it for its protection.
ansanolla@naic edu.Thursday, December 19, 2002 16:01:23Cryocomp was installed back into its shack . It fell from it.
ansanolla@naic.eduMonday, January 13, 2003 08:40:02System hot since 1/11/2003. Why ????, we don't know yet have to check.
ansanolla@naic edu.Tuesday, January 14, 2003 15:38:32The cryocomp was replaced ( was 012 , is 009 ).Going to be cooled tomorrow.
ansanolla@naic edu.Wednesday, January 15, 2003 11:26:51System been cooled down , cryocomp is 009
ansanolla@naic edu.Tuesday, January 21, 2003 14:41:40Cryo lines 19/20 were replaced with cryo lines 17/18.Vives is finihing lines 19/20.
ansanolla@naic edu.Thursday, January 30, 2003 08:56:21System down for modifications.RF sw's swap.
ans anolla@naic edu.Thursday, March 6, 2003 16:10:18System back to normal after rf sw replacement.
lswlwray@naic.eduThursday, March 6, 2003 16:35:48problem with cals not firing traced to cal box -- 15V supply to diodes not making connection. fixed today.
lswlwray@naic.eduWednesday, March 26, 2003 16:24:55removed the X Band from the SBH/XB combination.
ansanolla@naic.eduMonday, April 28, 2003 11:48:51Post amp chassis down in order to replace cal diodes.
ansanolla@naic.eduMonday, May 5, 2003 13:00:55Having problems with the new noise diodes .They Are not switching properly , they even fire with less than .2 vdc ttl drive.
ansanolla@naic.eduTuesday, May 6, 2003 12:56:00The SBH receiver was installed back.Was out since 4/28/2003 in order to replace the cal diodes , it was not posible due that the new cal diodes were defective. It's ttl drive level for low was .2 vdc .So the old cal diodes are back on it for a while. it's tsys was for chann A 47 k and 48 k for chann B.
ansanolla@naic.eduMonday, May 12, 2003 14:43:19The cryosystem is leaking He at a rate of 20 psi/day. So for the importance of the experiment that is going to be run next week , the cryocooling system (He lines and cryocomp) from the SBW was swap to this receiver & viceversa. Now SBH is with the SBW cryo cooling system and viceversa.
ansanolla@naic.eduMonday, June 16, 2003 14:11:53The cryocomp ED 012 is back with the system , the 20 psi He leak was localized in the flex cryo line at the top of the Dome ,it was the return line.
ansanolla@naic.eduTuesday, August 5, 2003 15:02:52The cryopump was replaced , it was ED 028 . Now is ED 021.
ansanolla@naic.eduWednesday, August 6, 2003 11:59:50The rf was turned on.Cryo temps are 16 stage is 14.7 and 70 stage is 52.4.
ansanolla@naic.eduWednesday, November 19, 2003 08:40:14System has been warm since 11/9/2003 due that it's cryocomp (ED 010)was used to replace the Trio recvr cryocomp ( ED 011). It strted spilling oil. Also Cheo had it's control boxes down at the lab for modification.
ansanolla@naic.eduWednesday, November 19, 2003 11:19:12System been cool down. There is not cryo temp monitor due that chassis are down at the cryo shop in order to be modified.
ansanolla@naic.eduThursday, November 20, 2003 13:52:24The system is back again. Was down due that it's cryocomp was used for the trio receivers , but is back again.
ansanolla@naic.eduMonday, November 24, 2003 12:41:25Cryopump (ED021) is noisy.It has a bad ball bearing at the cryopump yoke.
ansanolla@naic.eduTuesday, December 16, 2003 14:23:19System down ( bias & rf chassis down for modification)
ansanollaq@naic.eduWednesday, December 17, 2003 13:34:25System back to normal. Due to 430 mhz xmitter failure all Dmoe recvrs back to operation.
lswlwray@naic.eduMonday, February 2, 2004 15:55:13Leak in SBN/XB system has been avoided by moving the lines to old#13 (receiver room to transmitter room) and new #17 from transmitter room up to compressor area. Receiver is cooling.
ansanolla@naic.eduFriday, February 13, 2004 16:15:30System down for modification . Going to replace cal diodes.
ansanolla@naic.eduWednesday, March 3, 2004 10:09:28Cal noise diodes replaced with Noise Com ones. Are replacing Micronetics ones.
ansanolla@naic.eduWednesday, March 3, 2004 15:50:55System been cooled down. It's using cryocomp ED 011 at the gimbles with new cryo lines 17 (R) & 18 (S).New cal diodes added (Noise Com ones)
ansanolla@naic.eduTuesday, March 9, 2004 12:58:59New cal diodes were calibrated this morning by P. Perillat.
ansanolla@naic.eduMonday, May 3, 2004 16:20:51System down due to leaky cryo lines.
ansanolla@naic.eduTuesday, May 4, 2004 16:07:41flex cryo line 17 (R) at cable rap bad , it had a leak of about 20 psi/hour. Also line 18 (S) at rap had a leaky connector. System been cool.
ansanolla@naic.eduWednesday, May 5, 2004 12:54:13System back to normal after fixing cryo lines 17 & 18 .Both were leaking He , 17 had the flex line at cable rap bad & 18 had a cryo connector leaking at the end of cable rap.
lswlwray@naic.eduTuesday, June 1, 2004 16:03:52Tsys results for Friday May 29 showed odd results for the high cals but not the low cals. Also, saw very low levels for l90cal channel A when detecting the IF output. Problem linked to a bad inverter chip (U1) on the cal switching card in the postamp chassis. Chassis was taken down before lunch and put back at 3:30pm.
ansanolla@naic.eduFriday, July 2, 2004 11:31:59System down due to power shut down ( main power shorted at the rotary joint).
ansanolla@naic.eduSaturday, July 3, 2004 16:26:38System back to normal after cool down.
ansanolla@naic.eduMonday, October 11, 2004 15:56:57System was down due to power failure on Saturday , was cooled down by BG and JRV.
ansanolla@naic.eduThursday, October 14, 2004 14:32:12Cryo comp ED011 was leaking oil by the bunch , so it was replaced by cryo comp ED009. System been cooled down.
ansanolla@naic.eduFriday, October 15, 2004 12:48:11System cooled.
ansanolla@naic.eduWednesday, November 3, 2004 16:07:58System warming up.
ansanoll@naic.eduThursday, November 4, 2004 11:37:37There it was a vacuum leak on the Dewar thru the bias connector , was tight and vac seal on it.Also found cal 1 rf connector loose.
ansanolla@naic.eduTuesday, November 23, 2004 16:15:05Bias in channel A 1 was adjusted from 1.5 V & 5 ma to 1.4 V & 4.5 ma in order to kill an oscillation at 9.18 Ghz that was causing rfi to the XB receiver.
ansanolla@naic.eduTuesday, November 30, 2004 14:29:41The calibration system was not working properly in all the receivers at the Dome . It was due that the post amp chassis for the CBH and XB were off. These was loading the other recvrs calibrators.
lswlwray@naic.eduThursday, January 27, 2005 12:47:14Cheo and Nolla installed a new temp reading card with the new modifications on it (no 60 MHz comb). After this, we found a problem with reading temperatures on the Lakeshore, but reading remotely is fine, probably a wire that was disturbed while installing the new card.
ansanolla@naic.eduTuesday, February 22, 2005 16:24:28There it was a power shut down at the Dome due to a braker shut down at the substation that supplies the power to it.It affected several receivers because it last more than 30 minutes.The recvrs were 430 ,SBH,SBL and SBN.Vacuum pump on all of them.These happened on 2/18/2005
ansanolla@naic.eduTuesday, March 15, 2005 08:44:55Yesterday in the afternoon JRV noticed that the recvr. was warming up. So we track the problem to the cryocomp that shut off (Why?).Been pumped.
ansanolla@naic.eduFriday, April 1, 2005 10:49:13The cryo comp shut off due to a power dip that happens on 3/30/2005 at 22:30 hrs . So Coco started the cooldown yesterday (3/31/2005)in the morning time. System still have the vac valve open , have to be closed today.
ansanolla@naic.eduMonday, April 4, 2005 11:43:48System down for a while. Not needed for observation.Was turned off in order to use its cryo comp (009) replacing ALFA 2 (014)cryocomp that was bad (return pressure to high!!!!!!!
Bill Genterwgenter@naic.eduWednesday, April 13, 2005 16:52:57SBH Receiver is Operational at room temperature.
ansanolla@naic.eduMonday, May 9, 2005 13:36:03System been cooled , need to be used on 5/12/2005. Coco replaced it's cryohead , it was ED 029 , now is ED 030.Also cryocomp ED 009 was returned to system, it was with the ALFA (#2) receiver cryocool system.
ansanolla@naic.eduWednesday, May 11, 2005 11:53:35Receiver polarization test performed by BG.Only left polarization test was runned , no right side ant available.
ansanolla@naic.eduMonday, August 1, 2005 10:05:56Do to power dip on 7/29/2005 the cryo comp shut off. So Carlos Rios went upstairs to turn it on.Also he hooked up the pump into the Dewar cause it was warming up.Valve still open upstairs.
ansanolla@naic.eduFriday, August 26, 2005 13:06:11Due to power failure on 8/25/2005 system warmed up.
ansanolla@naic.eduTuesday, August 30, 2005 15:46:54System cooled. Back to normal operation.
ansanolla@naic.eduWednesday, August 31, 2005 16:09:33Shutter not working due to bad air compressor.Replace it.
ansanolla@naic.eduThursday, September 15, 2005 13:30:41The shutter set screw was set this time with Locktite.
ansanolla@naic.eduThursday, September 15, 2005 16:06:20For some reason system looks to be warming up.Carlitos & Calixto on way up to check problem.
crioscrios@naic.eduThursday, September 15, 2005 17:11:46SBH Temperature raise. Compresor 009 turned off after a power dip. We turned the Compresor on . SBH vaccum was almost at atmosphere.We connected a vaccum pump and turned the Crosshead when the vaccum was at 6 microns. SBH Temperature started to go down and the vaccum pump is still connected to the dewar.
ansanolla@naic.eduTuesday, October 18, 2005 12:55:14The shutter driver was switched to Ivan Cordero's shutter control box.CR worked on it.
ansanolla@naic.eduThursday, October 20, 2005 13:16:04A new safety divice added to the crosshead driver power sw in order to avoid to turn it off accidentally.
ansanolla@naic.eduWednesday, February 15, 2006 13:31:32SBH warmed up yesterday. Went upstairs to check and found that the Dewar pup up valve was out of place , like if a positive preassure was applied to the Dewar ,??????? No logic explanation to it.Been warmed up to remove contamination because it pick up a lot of it.Was tested for vacuum and was ok .First time it happen in one of our Dewars.
ansanolla@naic.eduFriday, February 17, 2006 15:23:36Vacuum Dewar was 20 un's , found that vacuum pump had a bad connector at the vac hose , replaced.Cool down will be started on Monday.
ansanolla@naic.eduMonday, February 20, 2006 11:32:48SBH cooldown started , Dewar vacuum was at 5 un's.
ansanolla@naic.eduTuesday, February 21, 2006 11:48:15System up again.Valve closed at the Dewar.
ansanolla@naic.eduWednesday, February 22, 2006 15:12:10Channel B first stage amp burned.Bias current won't adjust.
ansanolla@naic.eduThursday, February 23, 2006 12:39:39System down.Been warming up in order to replace the Dewar channel B amp, its first stage is burned out.
ansanolla@naic.eduTuesday, February 28, 2006 13:43:12Channel B amp replaced , it was burned out. The sn of the new one is sn 103.The one replaced is sn 101 , the model # of the amplifiers is C3.5-30H.
ansanolla@naic.eduTuesday, February 28, 2006 13:47:22I entered wrong data in the last entry. The sn of the replaced amp is sn 102.The sn 101 is the one on channel A.
ansanolla@naic.eduWednesday, March 1, 2006 13:22:20System been cooled down.Dewar vac was 3.5 un's.The recvr 16 deg stage monitor is not reading , has to find what is wrong with it.Also its cryo comp (ED 009) was modified by CR & WI for its time delay start protection , it was not wired for it.
ansanolla@naic.eduThursday, March 2, 2006 12:51:55The SBH system is cold.The 16 deg K monitor is not working properly something is loose inside the Dewar , so we are reading the 16 deg stage thru the OMT monitor due that its diode is on this stage.Bias for the the new amp on chann B (sn 103) are: Amp C-3.5-30H S/N 103 @ 15K Vd1 = 2.5V Id1= 8.5ma Vd2 = 3.0V Id2 = 8.0ma Vd3 = 3.0V Id3 = 10.0ma
ansanolla@naic.eduTuesday, August 15, 2006 13:46:23The SBH system cryo comp & lines were used to replace the ALFA 2 cryogenics due to a leak in these system. So SBH system is at room temp.
ansanolla@naic.eduThursday, August 17, 2006 09:08:35System cryogenics will be with the ALFA until new cryo hoses arrive for it.
ansanolla@naic.eduFriday, September 1, 2006 11:23:26The SBH Dewar being cooled after returning its cryo comp ED 009 to it.It was with the ALFA 2 cryo system since 8/17/2006.
ansanolla@naic.eduTuesday, September 5, 2006 15:28:51System back to normal after using its cryo system to cool down the ALFA 2 cryo system , it had a huge helium leak on it.
ansanolla@naic.eduThursday, October 5, 2006 12:52:50Cryo reading was not working properly.It was found that the monitor board on the bias chassis nor the output data cable were working properly both were bad.
crvcrios@naic.eduTuesday, December 12, 2006 13:18:34Cryo system has been shut off. Cryo comp. and lines being used for XB and SBN receivers.
ansanolla@naic.eduFriday, March 30, 2007 11:35:56System taken down for service.
ansanolla@naic.eduTuesday, May 1, 2007 11:21:38In the post amp chassis the power Sw & Lake Shore meter sw were replaced with lock type ones.
crvcrios@naic.eduTuesday, April 15, 2008 11:26:43Problem with Channel A calibration. channel a RF cable was broken . Cable was replaced with a new one that was phase matched ed with channel B cable.
crvTuesday, April 15, 2008 11:28:27The system is been vacuumed to be cooled tomorow.
crvFriday, April 18, 2008 11:16:34 The receiver had calibration problems. The amplifiers were off.
crvMonday, May 12, 2008 10:49:18This reciever is been warmed up.
crvFriday, June 6, 2008 10:43:26 On wednesday june the 4th on the afternoon due to problem on one of the phase voltage in the power company lines the compresor started cycling on and off. The dewar lost vacuum and started to warm up. We decided to let it warm completely and start vacuum on thursday. Today fraiday the 6th we started the cool down again.
crvWednesday, June 18, 2008 16:01:45the reciever has been warmed campaign end.
crvFriday, June 12, 2009 13:22:23System was cooled down on Wednesday june 10th 2009.
CrvWednesday, July 1, 2009 14:31:50Reciever warmed up on June 26th 2009 Campaign ended.
crvFriday, January 8, 2010 16:25:18System was vac. and is cooling down. System was connected to comp. ED015.
crvcriosThursday, March 4, 2010 13:36:36Monday, February 22, 2010 14:43:30 SBh was connected to compressor C2 of ALFA.
crvcrios@naic.eduThursday, March 11, 2010 13:36:48The system reached good cryo temp but it is racheting (contamination sound). The system is warming up with the contamination trapped on the Cryo pump to eliminate it from the system. Cryo pump will be purged and checked once the dewar reaches room temp.
CriosFriday, April 30, 2010 17:34:15The system warmed up due to comp. C2 failure on April 22nd.
crvWednesday, June 2, 2010 10:37:00The system was connected to compressor C3 on May 28th,2010 after removing 430 & CB from the compressor. Cooling will star on the weekend.
crvcrios@naic.eduWednesday, June 2, 2010 10:40:26The system is down agaian due to an internal broken line on the compressor. Below the main dish is still flooded and is dificult to lower or raise a compressor.
criosWednesday, May 25, 2011 15:30:50The receiver started cooldown on 5/13/2011.
Dana Whitlowdwhitlow@naic.eduWednesday, December 12, 2012 09:00:33The S-Band(high) receiver was taken out of service and brought down to the lab in late October 2012 due to a severe loss of performance on Pol A. Measurements in the lab quickly showed that LNA in Pol A was at fault, having suffered a decrease in gain to the tune of 20 dB from normal. We were unable to locate the spare unit that we supposedly had on hand. Next followed a series of exchanges with QuinStar Technologies, the company who bought Berkshire Technologies, the original maker of the LNA (model C-3.5-30H). This went on for weeks, culminating yesterday in an invitation for us to send the failed unit to QuinStar for evaluation. I had a voice call with Rosalind Shuler (RShuler@QUINSTAR.com; 310-891-2408). She indicated a high likelihood that they would be unable to effect repairs due to the unit's extreme age, in which case we would have to have custom amplifier(s) built from scratch. The amplifier is now packaged and expected to ship out today via FedEx.
Denis Urbaindurbain@naic.eduFriday, August 9, 2013 16:47:45The receiver had to be removed from the rotary floor to leave room for AO19. When it was re-installed the amplifiers could not be biased. Carlos replaced the bias board with a spare one and it fixed the problem. The Cal LED is not turning ON and OFF, we suspect that the LED is dead because the Tsys gives the good temperatures.
SBH is now upstarisfelix.fernandez@naic.eduTuesday, February 10, 2015 11:45:41SBH was brought up last week and today all the helium lines were connected to an independent compressor (M450). Vacuum is currently at 2mTorr and cold down should start this afternoon.
SBH back to operational statusfelix.fernandez@naic.eduWednesday, February 11, 2015 15:02:07SBH is back to operational status.
SBH new amplifiersfelix.fernandez@naic.eduWednesday, February 11, 2015 15:06:40During SBH's last repair work the LNA were replaced by a set built by NRAO. These LNA have 4 bias stage each (Vd=1V and ID=5mA). Thus, we rewired the whole inside to accommodate for this 8 bias stages. Furthermore, since these LNAs are InP the clamping diodes structures were redesigned (very crudly) according to NRAO advice.
SBH ready to usecrios@naic.eduThursday, April 9, 2015 09:59:13System is cold and ready to be used again.
Cryo system maintnace.crios@naic.eduTuesday, September 22, 2015 14:29:04The cryo system was racheting ocasionaly and as time passed it was more often . Cryo was turned off Yesterday to replace the cryopump that was on operation for more than a year. Cryo pump replaced and started cooldown today.
Compressor and Cold-head replacementfelix.fernandez@naic.edTuesday, October 27, 2015 11:09:47During the past week we replaced the cold-head thinking it had a mechanical issue. It turned out it was not the cold-head but the compressor. The compressor had a broken valve at piston #1 (close to exterior). We re-routed 430/CB compressor to extend SBH observations while also replacing SBH's cold-head. This new cold-head started ratcheting. We tried flushing the cold-head only to have the cold-head blow its fuse (located in the Scott-T box) once reconnected. We will now warm-up SBH, replace the cold-head, and refurbish its compressor.
Cold-head contaminationfelix.fernandez@naic.eduFriday, January 8, 2016 09:29:59On Jan-4-2016 the SBN and then SBH's cold-head started ratcheting, most likely due to contamination. We started the warm-up procedure and will replace the cold-head after the WD campaign.
Cold-head failurefelix.fernandez@naic.eduTuesday, May 9, 2017 08:19:44On May 6, 2017 the receiver started to warm-up without any "warning". The compressor was working properly but not cooling was achieved by the cold-head. We replaced the cold-head on May 8 and started the cooling process on the same day.
Cold-head failurefelix.fernandez@naic.eduFriday, January 19, 2018 11:24:19On Jan-17-2018 SBH's cold-head failed suddenly. After inspection we found out that there was a mechanical failure. The cold-head was replaced on Jan-18 and on Jan-19 the vacuum process was started. We hope to have the receiver back by Jan-22.
Pol-A's LNA is 5dB below Pol-Bfelix.fernandez@naic.eduTuesday, January 30, 2018 14:32:43SBH pol-A LNA is 5dB below pol-B. Today we went to the Dome and plugged the spectrum analyzer to the dewar's output. Pol-A showed proper pass-band gain but it was 5dB below Pol-B's. Pol-A's LNAs bias of stage 1 is drawing twice as much current. It cannot be tuned using the bias chassis. Most likely we have a blown gate on the LNA and needs to be replaced.
SBH brought down for repairsfelix.fernandez@naic.eduTuesday, February 6, 2018 08:11:27SBH was brought down yesterday from the Dome for repairs. Main task will be to asses and repair Pol-A LNA.
Thermal cycle and LNA issuesfelix.fernandez@naic.eduWednesday, March 4, 2020 14:11:15The SBH receiver lost its vacuum due to an unexpected He leak. The leak degraded the cryogenic performance and the receiver started loosing its vacuum. Also, the PolA LNA has been drawing to much current on stage 1. Furthermore, trying to adjust the LNA's stage current did not show any changes in current level. Hence, it is clear that the gate of the LNA is not controlling the current. The same issue is also present in stage 3 of PolB. This issue arose during the unexpected thermal cycle.
SBH LNA replacementfelix.fernandez@naic.eduThursday, July 9, 2020 13:18:44SBH had (pre-pandemic) been having erratic bias values. After bringing down the receiver we found out that both LNAs had no gate control in stage 1 & 2. During the repairs we also found that the spare LNA would lose gate control at stage 1 once the LNA's temperature would go below 256K. The final solution was to replace the LNAs by: PolA - Chalmers 2-4GHz S/N 2006-03, this is SBN spare LNA PolB - Berkshire C-3.5-30H S/N 103, original SBH LNA replacements There were some loose connections inside the connector at the BiasChasis - Dewar cable. We had to replace the semirigi sma cables between the LNA output and the dewar's RF port on the 20K plate. This was due to the size difference of the LNAs we used for replacements. The new cables are a bit longer and added about 0.7dB extra loss (at room temp). SBH is now back online and ready for observations.

Back to Top